# “Rosenbluth” separation of the $J/\psi$ near-threshold photoproduction - an access to the gluon Gravitational Form Factors at high $t$

### v2

@@ Abstract
We perform analysis of the near-threshold $J/\psi$ photoproduction data off the proton based on two theoretical approaches, GPD Guo et al. (2024) and holographic Mamo and Zahed (2022), that represent the differential cross sections as powers of the skewness parameter with coefficients that depend only on the momentum transfer $t$. This allows to separate kinematically the corresponding coefficient functions, in much the same way as this is done for the electric and magnetic form factors using the Rosenbluth separation. We examine the independence of the extracted functions with the photon beam energy. These functions, under additional assumptions, are related to the proton’s gluon Gravitational Form Factors (gGFFs). We compare the extracted functions with lattice calculations of the gGFFs in the region of $20.5<|t|<$ 2 GeV2, where they overlap. Such analysis demonstrates the possibility of extracting some combinations of the gGFFs from the data at high $t$, complementary to the lattice calculations available in the low $t$ region. However, higher statistics are needed to more accurately check the predicted scaling behavior of the data and compare with the lattice results, thus testing and comparing the theoretical assumptions used in the GPD and holographic models.
preprint:
@@ I Introduction
In Ref. Duran et al. (2023) (J/$\psi$-007 collaboration) the gGFFs have been extracted by fitting the $J/\psi$-007 data with the theoretical predictions from the holographic Mamo and Zahed (2022) and GPD Guo et al. (2021) approaches. In the GPD analysis of Ref. Guo et al. (2023), the data from both GlueX and $J/\psi$-007 experiments were used to extract the gGFFs. The GPD approach is based on expansion in the skewness parameter $\xi$ for values approaching unity (see also Hatta and Strikman (2021)). Therefore its validity is justified for high $|t|$ values due to kinematic correlation between $t$ and $\xi$. A common feature of the above two analyses in Refs. Duran et al. (2023) (J/$\psi$-007 collaboration); Guo et al. (2023) is the modeling of the gGFFs with dipole/tripole functions. The dipole form was proposed in Ref. Frankfurt and Strikman (2002) in analogy with the electro-magnetic form factors, whereas the tripole one is a good approximation of the holographic results Mamo and Zahed (2022). Such functional forms are consistent with the lattice calculations in Refs. Pefkou et al. (2022); Hackett et al. (2023) and used there as a parametrization of the results. While the GPD approach works at high $|t|$, the lattice calculations of the gGFFs have been done so far in the $20<|t|<$ 2 GeV2 region. Therefore, the use of dipole/tripole parametrization especially at high $|t|$ requires experimental validation. Most importantly, the extraction of the gGFFs in both analysis of Refs. Duran et al. (2023) (J/$\psi$-007 collaboration); Guo et al. (2023) is guided by additional external constraints on some of the gGFF parameters, based on, or consistent with the lattice calculations. As a result, first, the extracted gGFFs are sensitive to the normalization of the theoretical models and second, when compared to the lattice calculations, such results are not completely independent.
@@
making direct connection to the gravitational form factors $A_{g}(t)$ and $C_{g}(t)$, where $\eta$ has a similar meaning as the skewness parameter $\xi$. It is important to stress that the GPD and holographic theories have very different domains of validity. The holographic theory is valid for soft processes because it works in the limit of high coupling constant. In contrast the GPD theory requires high $|t|$ and works for hard processes. We aim to extract the $t$-functions in Eqs.(1,2) (we will call them form factor functions) from the experimental data without any additional assumptions. In contrast to the studies in Refs. Duran et al. (2023) (J/$\psi$-007 collaboration); Guo et al. (2023), first, the procedure described in this work does not use any external constraints and second, the results will not be model functions fitted to the data, but actual data points measuring the absolute value and $t$-dependence of the above form factors.
@@
where $\left(\frac{d\sigma}{d\Omega}\right)_{M}$ is the Mott cross section, $\tau=-t/4m^{2}$, where $m$ is the proton mass. By measuring $\sigma_{R}$ at the same momentum transfer $t$ as function of $\epsilon/\tau$, one can estimate the slope and the intercept, being respectively the electric and magnetic form factors squared at this value of $t$.
@@ II The “Rosenbluth” technique
$G_{2}(t)$ $=2{\cal{A}}_{g}(t){\cal{C}}_{g}(t)+2\frac{t}{4m^{2}}{\cal{B}}_{g}(t){\cal{C}}_{g}(t)-\left({\cal{A}}_{g}(t)+{\cal{B}}_{g}(t)\right)^{2}$ (7)
@@
In the leading-moment approximation the functions ${\cal{A}}_{g}$, ${\cal{B}}_{g}$, and ${\cal{C}}_{g}$ are related to the gluonic Gravitational Form Factors (gGFF) $A_{g}$, $B_{g}$, and $C_{g}$ as:
@@
$\eta$ $=\frac{M^{2}}{2(s-m^{2})-M^{2}+t},$ (13)
In this approach the $B_{g}$ form factor is zero. The $N(E)$ function is normalized using GlueX total cross section, which is in contrast to the GPD calculations that are absolute. The holographic theory is valid in the double limit of strong gauge coupling and large number of colors, $N_{c}$. To indicate the approximate nature of the results we will rewrite the above formula in analogy with Eq.(4) as:
@@
In the GPD case we will renormalize the $G$ functions in Eq.(4) as:
$\frac{d\sigma}{dt}(\xi,t)=F(E)\cdot 4(A_{1}^{conf})^{2}\xi^{-4}\left[G_{0}(t)+\xi^{2}G_{2}(t)+\xi^{4}G_{4}(t)\right],$ (18)
so that in the leading-moment approximation, Eqs.(6-11), when neglecting $B_{g}$, the $G_{0}$ and $H_{0}$ function have the same normalization, in which case we have similar relations to the gGFFs:
@@
Thus, in the further formulas we will attach the factor $4(A_{1}^{conf})^{2}$ to the $F(E)$ function.
@@
$\sigma_{R0}(E,t)=\frac{d\sigma}{dt}(E,t)\frac{\xi^{2}(E,t)}{F(E)}=\xi^{-2}(E,t)G_{0}(t)+G_{2}(t)+\xi^{2}(E,t)G_{4}(t),$ (22)
@@
Now the last term can be neglected with respect to the first one, and we have the $G_{0}$ function extracted from experimental data:
$G_{0}^{exp.}(t)=\left[\sigma_{R0}(E_{i},t)-\sigma_{R0}(E_{j},t)\right]/\left[\xi^{-2}(E_{i},t)-\xi^{-2}(E_{j},t)\right]$ (24)
@@
$(G_{0}(t)+G_{2}(t))^{exp.}=\left[\sigma_{R02}(E_{i},t)-\sigma_{R02}(E_{j},t)\right]/\left[\frac{\xi^{2}(E_{i},t)}{1-\xi^{2}(E_{i},t)}-\frac{\xi^{2}(E_{j},t)}{1-\xi^{2}(E_{j},t)}\right],$ (25)
@@ III.1 The GPD approach
Figure 2: Global five-parameter fit of the reduced cross sections, $\xi^{2}/F\cdot d\sigma/dt$, as function of $|t|$ (in GeV2), using the GPD Eq.(18) and the parametrizations given by Eqs. (27,28), from GlueX (red) and $J/\psi$-007 (black) experiments for different photon energies as indicated. The data points with $\xi<0.4$ are plotted (blue) but not used in the fit. For the further analysis only the fitted function at $10.82$ GeV is used.
In principle, we can fit only the reference cross section with some function, so that we can use it in Eqs.(24) and (25) to subtract the reduced cross sections at the $t$ values of the data points at the other energies. Still, as seen in Fig.1, due to the $\xi>0.4$ constraint, the $1-2$ GeV2 region in $|t|$ is not covered by the reference cross section, therefore it requires extrapolation to lower $|t|$. For that we will fit all the cross sections simultaneously using Eq. (18). Thus, the extrapolation of the reference energy will be guided by the data at the other energies, at the same time reducing the error of the function fitted to the reference energy.
@@
$G_{0}(t)$ $=\frac{G_{0}(0)}{(1+t/m_{G0})^{4}}$ (27)
$G_{2}(t)$ $=\frac{G_{2}(0)}{(1+t/m_{G2})^{4}}+conts_{G2},$ (28)
where the constant term added for $G_{2}$ improves the fit quality for $|t|>5$ GeV2. The fit results are shown in Fig.2. The data points for $\xi<0.4$, also plotted in the figure but not used in the fit, deviate from the fitted functions. In the further analysis for the extraction of $G_{0}^{exp.}(t)$, in Eq.(24) we will use only the fitted function for the reference energy $E_{i}=10.82$ GeV (Fig.2 bottom right).
@@ III.2 Holographic approach
Figure 5: Global five-parameter fit of the reduced cross sections, $\eta^{-2}/N\cdot d\sigma/dt$, as function of $|t|$ (in GeV2), using the holographic Eq.(14), from GlueX (red) and $J/\psi$-007 (black) experiments for different photon energies as indicated. For the further analysis only the fitted function at $10.82$ GeV is used.
@@
$H_{0}^{exp.}(t)=\left[\sigma_{H0}(E_{i},t)-\sigma_{H0}(E_{j},t)\right]/\left[\eta^{-2}(E_{i},t)-\eta^{-2}(E_{j},t)\right],$ (29)
@@
Following the same procedure as in the case of the GPD analysis, we perform global fit of the reduced cross sections using Eq. (14), where $H_{0}$ and $H_{2}$ functions are parametrized in the same way as $G_{0}$ and $G_{2}$ in Eqs. (27,28) and $H_{4}$ is neglected. The most important difference is that in the holographic case all the data points with $E>9.3$ GeV are included in the fit and there is no need to extrapolate the reference cross section at $E_{i}=10.82$ GeV as its $t$-range covers the range of all the other measurements, see Fig. 1. The fit results shown in Fig. 5 demonstrate good fit quality.
@@
$H_{2}^{exp.}(t)=\left[\sigma_{H2}(E_{i},t)-\sigma_{H2}(E_{j},t)\right]/\left[\eta^{2}(E_{i},t)-\eta^{2}(E_{j},t)\right],$ (31)
@@
The results for the extracted $H_{0}$ and $H_{2}$ form factor functions are shown in Fig. 6. They are fitted with squared dipole functions, demonstrating that in the case of the holographic calculations, the extracted data points are also energy independent. As there is no constraint on the skewness parameter, the extracted form factor functions extend down to lower $|t|$ values, as compared to the GPD case. Note, however, that in the low $|t|$ region the errors are much bigger than in the GPD analysis, as the errors of the measurements are scaled by bigger factors in Eqs.(30,32), compared to Eqs.(22,26), in the corresponding reduced cross sections.
@@ IV Relation to the gluon Gravitational Form Factors
In Fig. 7(a) we compare the functions fitted to the data points (as extracted in the previous section) for $G_{0}$ and $H_{0}$. We see a general agreement between the GPD and holographic results, which is remarkable as these two theories are diametrically different. A possible way to interpret their agreement is that the corrections to both theories, which have very different nature, are not dominant. Indeed, according to Eqs.(19,15), the leading terms of both, $G_{0}$ and $H_{0}$, are the same and equal to the $A_{g}$ form factor squared. This assumption is supported by comparing the experimental results to the lattice calculations of $A_{g}^{2}$ from Ref. Hackett et al. (2023) (shown in the same figure), where we see an agreement at least within $2\sigma$ in the common $0.5-2$ GeV2 region.
In case of the $G_{0}+G_{2}$ and $H_{0}$ their leading terms should be proportional to the product of the gGFFs $A_{g}C_{g}$ (Eqs.(20,16)). The functions fitted to data points for $(G_{0}+G_{2})/8$ and $H_{0}/8$ from the previous section are shown in Fig. 7(b) and compared to the lattice calculation of $A_{g}C_{g}$. The agreement is not as good as in the $A_{g}^{2}$ case, at the same time the uncertainties of the data are much bigger so that we cannot make a definitive conclusion.
@@
We will discuss the vanishing of the terms containing $G_{4}$ and $H_{4}$ in Eqs.(18,14) within the leading-term approximation used in this section. Note that the mass scale parameter for $G_{0}$ is bigger than the one for $G_{0}+G_{2}$, $m_{G0}>m_{G02}$ (Fig. 4). The same is true when comparing the results for $H_{0}$ and $H_{2}$: $m_{H0}>m_{H2}$ (Fig. 6). According to Eqs.(19,20) and Eqs.(15,16), this means that $C_{g}$ vanishes faster with $|t|$ than $A_{g}$. As $G_{4}$ and $H_{4}$ are proportional to $C_{g}^{2}$ (Eqs.(21,17)), this explains qualitatively their vanishing contribution in Eqs.(18,14) at high $|t|$ values. At low $|t|$ where $\xi$ and $\eta$ are much smaller than unity, these terms are suppressed by the $\xi^{4}$/$\xi^{2}$, or $\eta^{4}$/$\eta^{2}$ factors when compared to the first/second term in Eqs.(18,14). From the extracted $A_{g}^{2}$ and $A_{g}C_{g}$ combinations we can evaluate separately the $C_{g}$ form factor and then calculate the $G_{4}$ and $H_{4}$ functions using Eqs.(21,17). In this way we have verified quantitatively, based on the results in this paper, that for $|t|>0.7$ GeV2, $\xi^{4}G_{4}$ is much smaller in absolute value than $G_{0}$ and $\xi^{2}G_{2}$. The same is true in the holographic case: $\eta^{4}H_{4}$ is much smaller in absolute value than $H_{0}$ and $\eta^{2}H_{2}$ for $|t|>0.5$ GeV2. Thus, in the leading-term approximation, this justifies neglecting the $G_{4}$/$H_{4}$ terms when extracting the form factors.

### v3

@@ Abstract
We perform analysis of the near-threshold $J/\psi$ photoproduction data off the proton based on two theoretical approaches, GPD Guo et al. (2024) and holographic Mamo and Zahed (2022), that represent the differential cross sections as powers of the skewness parameter with coefficients that depend only on the momentum transfer $t$. This allows to separate kinematically the corresponding coefficient functions in much the same way as this is done for the electric and magnetic form factors using the Rosenbluth separation. We examine the independence of the extracted functions of the photon beam energy. These functions, under additional assumptions, are related to the proton’s gluon Gravitational Form Factors (gGFFs). We compare the extracted functions with lattice calculations of the gGFFs in the region of $20.5<|t|<$ 2 GeV2, where they overlap. Such analysis demonstrates the possibility of extracting some combinations of the gGFFs from the data at high $t$, complementary to the lattice calculations available in the low $t$ region. However, higher statistics are needed to more accurately check the predicted scaling behavior of the data and compare with the lattice results, thus testing and comparing the theoretical assumptions used in the GPD and holographic models.
@@ I Introduction
In Ref. Duran et al. (2023) (J/$\psi$-007 collaboration) the gGFFs have been extracted by fitting the $J/\psi$-007 data with the theoretical predictions from the holographic Mamo and Zahed (2022) and GPD Guo et al. (2021) approaches. In the GPD analysis of Ref. Guo et al. (2023), the data from both GlueX and $J/\psi$-007 experiments were used to extract the gGFFs. The GPD approach is based on expansion in the skewness parameter $\xi$ for values approaching unity (see also Hatta and Strikman (2021)). Therefore, its validity is justified for high $|t|$ values due to kinematic correlation between $t$ and $\xi$. A common feature of the above two analyses in Refs. Duran et al. (2023) (J/$\psi$-007 collaboration); Guo et al. (2023) is the modeling of the gGFFs with dipole/tripole functions. The dipole form was proposed in Ref. Frankfurt and Strikman (2002) in analogy with the electro-magnetic form factors, whereas the tripole one is a good approximation of the holographic results Mamo and Zahed (2022). Such functional forms are consistent with the lattice calculations in Refs. Pefkou et al. (2022); Hackett et al. (2023) and used there as a parametrization of the results. While the GPD approach works at high $|t|$, the lattice calculations of the gGFFs have been done so far in the $20<|t|<$ 2 GeV2 region. Therefore, the use of dipole/tripole parametrization, especially at high $|t|$ requires experimental validation. Most importantly, the extraction of the gGFFs in both analyses of Refs. Duran et al. (2023) (J/$\psi$-007 collaboration); Guo et al. (2023) is guided by additional external constraints on some of the gGFF parameters, based on, or consistent with the lattice calculations. As a result, first, the extracted gGFFs are sensitive to the normalization of the theoretical models and second, when compared to the lattice calculations, such results are not completely independent.
@@
making direct connection to the gravitational form factors $A_{g}(t)$ and $C_{g}(t)$, where $\eta$ has a similar meaning as the skewness parameter $\xi$. It is important to stress that the GPD and holographic theories have very different domains of validity. The holographic theory is valid for soft processes because it works in the limit of high coupling constant. In contrast, the GPD theory requires high $|t|$ and works for hard processes. We aim to extract the $t$-functions in Eqs.(1,2) (we will call them form factor functions) from the experimental data without any additional assumptions. In contrast to the studies in Refs. Duran et al. (2023) (J/$\psi$-007 collaboration); Guo et al. (2023), first, the procedure described in this work does not use any external constraints and second, the results will not be model functions fitted to the data, but actual data points measuring the $t$-dependence of the above form factors.
@@
where $\left(\frac{d\sigma}{d\Omega}\right)_{M}$ is the Mott cross section, $\tau=-t/4m^{2}$, $\epsilon=1/(1+2(1+\tau)tan^{2}\frac{\theta_{e}}{2})$, where $m$ is the proton mass and $\theta_{e}$ is the electron scattering angle in the proton rest frame. By measuring $\sigma_{R}$ at the same momentum transfer $t$ as function of $\epsilon/\tau$, one can estimate the slope and the intercept, being respectively the electric and magnetic form factors squared at this value of $t$.
@@ II The “Rosenbluth” technique
$G_{2}(t)$ $=2{\cal{A}}_{g}(t){\cal{C}}_{g}(t)+2\frac{t}{4m^{2}}{\cal{B}}_{g}(t){\cal{C}}_{g}(t)$ (7)
$-\left({\cal{A}}_{g}(t)+{\cal{B}}_{g}(t)\right)^{2}$
@@
In the leading-moment approximation the functions ${\cal{A}}_{g}$, ${\cal{B}}_{g}$, and ${\cal{C}}_{g}$ are related to the gGFFs, $A_{g}$, $B_{g}$, and $C_{g}$ as:
@@
$\eta$ $=\frac{M^{2}}{2(s-m^{2})-M^{2}+t}.$ (13)
In this approach the $B_{g}$ form factor is zero. The $N(E)$ function as given in Ref. Mamo and Zahed (2022) depends only on the beam energy [^1] In the previous work of Ref. Mamo and Zahed (2020) this prefactor function in the Eq.(8.6) has a weak dependence on $t$. As it varies no more than $10\%$ over the full kinematic region considered here, the $t$ dependence will be neglected. and is normalized using GlueX total cross section. This is in contrast to the GPD calculations that are absolute. The holographic theory is valid in the double limit of strong gauge coupling and large number of colors, $N_{c}$. To indicate the approximate nature of the results we will rewrite the above formula in analogy with Eq.(4) as:
@@
In the GPD case we can rewrite Eq.(4) as:
$\frac{d\sigma}{dt}(\xi,t)$ $=F(E)\cdot 4(A_{1}^{conf})^{2}\xi^{-4}$ (18)
×$\times\left[G_{0}(t)+\xi^{2}G_{2}(t)+\xi^{4}G_{4}(t)\right].$
Thus, in the leading-moment approximation and neglecting $B_{g}$, from Eqs.(6-11) we obtain for the $G$ functions the same normalization and similar relations to the gGFFs as those for the $H$ functions:
@@
In the further formulas we will attach the factor $4(A_{1}^{conf})^{2}$ to the $F(E)$ function.
@@
$\sigma_{R0}$ $(E,t)=\frac{d\sigma}{dt}(E,t)\frac{\xi^{2}(E,t)}{F(E)}=$ (22)
$=\xi^{-2}(E,t)G_{0}(t)+G_{2}(t)+\xi^{2}(E,t)G_{4}(t),$
@@
The last term can be neglected with respect to the first one, and we have the $G_{0}$ function extracted from experimental data:
$G_{0}^{exp.}(t)$ $=\left[\sigma_{R0}(E_{i},t)-\sigma_{R0}(E_{j},t)\right]$ (24)
$/\left[\xi^{-2}(E_{i},t)-\xi^{-2}(E_{j},t)\right]$
@@
$(G_{0}(t)$ $+G_{2}(t))^{exp.}=\left[\sigma_{R02}(E_{i},t)-\sigma_{R02}(E_{j},t)\right]$ (25)
$/\left[\frac{\xi^{2}(E_{i},t)}{1-\xi^{2}(E_{i},t)}-\frac{\xi^{2}(E_{j},t)}{1-\xi^{2}(E_{j},t)}\right],$
@@ III.1 The GPD approach
Figure 2: Global five-parameter fit of the reduced cross sections, $\xi^{2}/F\cdot d\sigma/dt$, as function of $|t|$ (in GeV2), using the GPD Eq.(18) and the parametrizations given by Eqs. (27,28), from GlueX (solid red at $9.86$ and $10.82$ GeV) and $J/\psi$-007 (solid black) experiments for different photon energies as indicated. The data points with $\xi<0.4$ are plotted (open blue) but not used in the fit. For the further analysis only the fitted function at $10.82$ GeV is used.
In principle, we can fit only the reference cross section with some function of $t$, so that we can use it in Eqs.(24) and (25) to subtract the reduced cross sections at the $t$ values of the data points at the other energies. Still, as seen in Fig.1, due to the $\xi>0.4$ constraint, the $1-2$ GeV2 region in $|t|$ is not covered by the reference cross section, therefore it requires extrapolation to lower $|t|$. For that we will fit all the cross sections simultaneously using Eq. (18). Thus, the extrapolation of the reference energy will be guided by the data at the other energies, at the same time reducing the error of the function fitted to the reference energy.
@@
$G_{0}(t)$ $=\frac{G_{0}(0)}{(1-t/m_{G0}^{2})^{4}}$ (27)
$G_{2}(t)$ $=\frac{G_{2}(0)}{(1-t/m_{G2}^{2})^{4}}+conts_{G2}.$ (28)
The constant term added for $G_{2}$ improves the fit quality for $|t|>5$ GeV2 and it is not important for the extrapolation of the reference cross section to lower $|t|$ values. The fit results are shown in Fig.2. The data points for $\xi<0.4$, also plotted in the figure but not used in the fit, deviate from the fitted functions. In the further analysis for the extraction of $G_{0}^{exp.}(t)$, in Eq.(24) we will use only the fitted function for the reference energy $E_{i}=10.82$ GeV (Fig.2 bottom right).
@@ III.2 Holographic approach
Figure 5: Global five-parameter fit of the reduced cross sections, $\eta^{-2}/N\cdot d\sigma/dt$, as function of $|t|$ (in GeV2), using the holographic Eq.(14), from GlueX (solid red at $9.86$ and $10.82$ Ge) and $J/\psi$-007 (solid black) experiments for different photon energies as indicated. For the further analysis only the fitted function at $10.82$ GeV is used.
@@
$H_{0}^{exp.}(t)$ $=\left[\sigma_{H0}(E_{i},t)-\sigma_{H0}(E_{j},t)\right]$ (29)
$/\left[\eta^{-2}(E_{i},t)-\eta^{-2}(E_{j},t)\right],$
@@
Following the same procedure as in the case of the GPD analysis, we perform global fit of the reduced cross sections using Eq. (14), where $H_{0}$ and $H_{2}$ functions are parametrized in the same way as $G_{0}$ and $G_{2}$ in Eqs. (27,28) and $H_{4}$ is neglected. The most important difference is that in the holographic case all the data points with $E>9.3$ GeV are included in the fit and there is no need to extrapolate the reference cross section at $E_{i}=10.82$ GeV as its $t$-range covers the ranges of all the other measurements, see Fig. 1. The fit results shown in Fig. 5 demonstrate good fit quality.
@@
$H_{2}^{exp.}(t)$ $=\left[\sigma_{H2}(E_{i},t)-\sigma_{H2}(E_{j},t)\right]$ (31)
$/\left[\eta^{2}(E_{i},t)-\eta^{2}(E_{j},t)\right],$
@@
The results for the extracted $H_{0}$ and $H_{2}$ form factor functions are shown in Fig. 6. They are fitted with squared dipole functions, demonstrating that in the case of the holographic calculations, the extracted data points are also energy independent. As there is no constraint on the skewness parameter, the extracted form factor functions extend down to lower $|t|$ values, as compared to the GPD case. Note, however, that especially in the low $|t|$ region the errors are much bigger than in the GPD analysis, as the errors of the measurements are scaled by bigger factors in Eqs.(30,32), compared to Eqs.(22,26), in the corresponding reduced cross sections.
@@ IV Relation to the gluon Gravitational Form Factors
In Fig. 7(a) we compare the functions fitted to the data points (as extracted in the previous section) for $G_{0}$ and $H_{0}$. We see a general agreement between the GPD and holographic results, which is remarkable as these two theories are diametrically different. A possible way to interpret their agreement is that the corrections that have very different nature for the two theories, are not dominant. Indeed, according to Eqs.(19,15), the leading terms of both, $G_{0}$ and $H_{0}$, are the same and equal to the $A_{g}$ form factor squared. This assumption is supported by comparing the experimental results to the lattice calculations of $A_{g}^{2}$ from Ref. Hackett et al. (2023) (shown in the same figure), where we see an agreement at least within $2\sigma$ in the common $0.5-2$ GeV2 region.
The leading terms of $G_{0}+G_{2}$ and $H_{2}$ should be proportional to the product of the gGFFs $A_{g}C_{g}$ (Eqs.(20,16)). The functions fitted to data points for $(G_{0}+G_{2})/8$ and $H_{2}/8$ from the previous section are shown in Fig. 7(b) and compared to the lattice calculation of $A_{g}C_{g}$. The agreement is not as good as in the $A_{g}^{2}$ case, at the same time the uncertainties of the data are much bigger so that we cannot make a definitive conclusion.
@@
We will discuss the vanishing of the terms containing $G_{4}$ and $H_{4}$ in Eqs.(18,14) within the leading-term approximation used in this section. Note that the mass scale parameter for $G_{0}$ is bigger than the one for $G_{0}+G_{2}$, $m_{G0}>m_{G02}$ (Fig. 4). The same is true when comparing the results for $H_{0}$ and $H_{2}$: $m_{H0}>m_{H2}$ (Fig. 6). According to Eqs.(19,20) and Eqs.(15,16), this means that $C_{g}$ vanishes faster with $|t|$ than $A_{g}$. As $G_{4}$ and $H_{4}$ are proportional to $C_{g}^{2}$ (Eqs.(21,17)), this explains qualitatively their vanishing contribution in Eqs.(18,14) at high $|t|$ values. At low $|t|$ where $\xi$ and $\eta$ are much smaller than unity, these terms are suppressed by the $\xi^{4}$/$\xi^{2}$, or $\eta^{4}$/$\eta^{2}$ factors when compared to the first/second term in Eqs.(18,14). From the extracted $A_{g}^{2}$ and $A_{g}C_{g}$ combinations we can evaluate separately the $C_{g}$ form factor and then calculate the $G_{4}$ and $H_{4}$ functions using Eqs.(21,17). In this way we have verified quantitatively, based on the results in this paper, that for $|t|>0.7$ GeV2, $\xi^{4}G_{4}$ is about an order of magnitude smaller in absolute value than $G_{0}$ and $\xi^{2}G_{2}$. The same is true in the holographic case: $\eta^{4}H_{4}$ is much smaller in absolute value than $H_{0}$ and $\eta^{2}H_{2}$ for $|t|>0.5$ GeV2. Thus, in the leading-term approximation, this justifies neglecting the $G_{4}$/$H_{4}$ terms when extracting the form factors.